A music store is selling compact disks for $11.50 and $7.50. You buy
12 disks and spend a total of $110. How many compact disks that cost
$11.50 did you buy?

Found 2 solutions by palanisamy, richwmiller:
Answer by palanisamy(496)

A music store is selling compact disks for $11.50 and $7.50.
Let the number of disks of first type bought = x
and the number of disks of second type bought = y
Total number of disks x+y = 12 ...(1)
Total amount 11.50x+7.50y = 110 ...(2)
(2)*2 => 23x+15y = 220 ...(3)
(1)*15=> 15x+15y = 180 ...(4)
(3)-(4)=> 8x = 40
x = 40/8
x = 5
(1) => 5+y = 12
y = 12-5
y = 7
Therefore the number of diks of price $ 11.50 = 5

Answer by richwmiller(17219)

s+e=12
s=12-e
s*7.50+e*11.50=110
(12-e)*7.50+e*11.50=110
Solve for e